# thanks: Dedicated to Prof. Jafar Sadeghi in honor of his 60th jubilee

### v1

@@ IV The new agegraphic dark energy model
where $\dot{\tau}=\frac{1}{a}$. Therefore, the energy density of the new agegraphic dark energy reads
@@
where $n^{2}$ is constant. In the next section, we will explore BEC model in the context of cosmological evolution.
V BEC model as dark matter
In this section, we intend to study the nature of dark matter as one of the dark universe parts by the regime of BEC. BEC is a state of matter in which the dilute Bose gas is cooled to such a low temperature that a phase transition occurs at this point and the macroscopic quantum phenomenon occurs as a quantum minimum state. Since dark matter has a significant contribution to the evolution of the universe, it leads us to adopt the BEC approach instead of dark matter. This means that dark matter is a type of BEC boson gas that dominates a part of the universe. Therefore,we consider the number density of particles by the Bose-Einstein statistics, which these particles are formed by decoupling of the remaining plasma in the early universe. However, the energy density of dark matter is equal to the mass of dark matter multiplied by the number density of particles, and dark matter pressure is defined by Bose-Einstein statistics in a sphere with a radius of particles momentum (refer to [89, 90, 91, 92, 93] for more details). In that case, by the generalized Gross-Pitaevskii equation [94, 95], we write down the dark matter pressure in terms of the energy density of dark matter as
@@ V BEC model as dark matter
where $\omega_{dm}=2\pi l_{a}/m_{dm}^{3}$, in which $l_{a}$ is the s-wave scattering length and here is introduced as EoS of dark matter. By inserting Eq. (20) into Eq. (13b), we have
@@
where by inserting $a=a_{0}=1$ ($a_{0}$ is the current scale factor) we obtain the current $\rho_{dm_{0}}$ as
@@
VI Reconstructing and correspondence of the model
In this section, we study the present model based on reconstructing and correspondence methods. In that case, we reconstruct the Friedmann equation in terms of redshift parameter, and then we correspond the agegraphic dark energy with $f(T)$ gravity. For this purpose, we write down the transformation between the scale factor and redshift parameter $z$ as $a(t)=a_{0}/(1+z)$ in which $a_{0}=1$ is the present scale factor, also differential form is as
@@ VI Reconstructing and correspondence of the model
where $t_{0}$ is the present age of the universe, and $m$ is dimensionless positive coefficient. The Hubble’s parameter and consequently the age of the universe will be:
@@
The interesting point is that we can calculate the value $m$ by fitting Eq. (27) with the astronomical data. For this purpose, we adopt the value of the coefficient $m$ from the result of Ref. [44] as $m=0.956$, from which we can immediately obtain the present age of the universe from Eq. (26b) as $t_{0}=13.75Gyr$. In what follows, we consider that the source of dark energy arises from the $f(T)$ gravity and the new agegraphic. Therefore, we correspond Eq. (19) with energy density of dark energy $\rho_{de}$ and obtains as
@@
Now, by entering equations (14), (23), (27), and (28)in Eq. (15a), we can obtain the form of the function $f(T(z))$ including the value of $b=0.807$ as follows:
$f(z)=53.068n^{2}(1+z)^{-1.138}-\frac{1.395\kappa^{2}\rho_{bm_{0}}}{2\omega_{bm}+1.303}(1+z)^{3\omega_{bm}+1.954}$
$+\kappa^{2}\rho_{dm_{0}}\left(\omega_{dm}-0.349\right)\left(11.737\ln\!\left(1+z\right)-5.735\ln\!\left(-0.349(1+z)+\omega_{dm}(1+z)^{2.046}\right)\right),$ (30)
where by using Eqs. (7) and (27), which results in $z=-1+\left(-T/H_{0}^{2}\right)^{n/2}$, we convert $f(z)$ as the form of the function $f(T)$ (it is unnecessary to write it here). In what follows, we can obtain the dark energy pressure $p_{de}$ by inserting Eq. (28) into the continuity equation (13c), and then we obtain the varition of $\rho_{de}$, $p_{de}$, and $\omega_{de}$ for dark energy component in terms of redshift parameter as shown in Figs. 1 and 2. We note that the free parameters has the key role for dynamical description of the cosmological parameters. In that case, we choice them based on positivity the dark energy density and negativity the dark energy pressure for the late universe ($z=0$). Therefore, the free parameters are selected as $n=18$, $\omega_{dm}=2.5$, $\rho_{dm_{0}}=4873$, $\omega_{bm}=1$, and $\rho_{bm_{0}}=545$ which $n$, $\omega_{dm}$, and $\omega_{bm}$ are dimensionless quantities, and $\rho_{dm_{0}}$ and $\rho_{bm_{0}}$ are dimension $M^{4}$ in Planck units. We can see in Figs. 1 that the value of energy density decrease from a high positive value in the early universe to a lower positive value in the late universe ($z=0$), also the value of pressure decrease from a high positive value in the early universe to a lower negative value in the late universe ($z=0$).
@@
We note that the EoS parameter is a very important parameter so that can understand at a glance how the dynamics of the universe are evolving from the early time to the late time. Therefore, we notice that the matter period and the accelerated pase period for the universe can be expressed as $0<\omega<\frac{1}{3}$ and $-1<\omega<-\frac{1}{3}$, respectively. Hence, Fig. 2 shows us that the universe begins from matter era to accelerated phase era in which the value of EoS is about $-0.71$ in the present time ($z=0$), that represent the universe is undergoing an accelerated expansion.
@@
In what follows, we explore the stability of the model by using the sound speed parameter which one comes from the concept of thermodynamics. For this reason, we consider the universe to be an adiabatic thermodynamic system where no heat or mass is transferred out, meaning that the perturbation has zero entropy. Therefore, the sound speed parameter $c_{s}^{2}$ is defined in the following form
@@
where the index $z$ indicates derivative with respect to redshift parameter. We now draw $c_{s}^{2}$ in terms of the redshift parameter as shown in Fig. 3. We note that the conditions of larger and smaller sound speed parameters, i.e., $c_{s}^{2}>0$ and $c_{s}^{2}<0$, indicate stability and instability, respectively. In that case, Fig. 3 displays that the universe begins from a stability phase and then at the threshold of the accelerated phase enters a non-stability phase in late time. Therefore, this variation represents that the energy density of dark energy in the late universe is not in a controlled growth.
Figure 3: The graph of the sound speed (dimensional quantity) in terms of redshift parameter by $\alpha=50$, $\beta=0.001$, $b=0.01$, $c=2$, $m=0.95$, $r=0.5$, $g=-1$, $\lambda=0.5$, $a_{0}=0.1$, and $\xi=0.25,1.25,2.25$.
@@
VII Conclusion
In this paper, we explored the universe in the context of $f(T)$ gravity by flat-FRW background. We considered the universe to be a combination of three components called baryonic matter, dark matter and dark energy. For this purpose, we took the origin of dark matter from BEC and considered the source of dark energy from the new agegraphic. After calculating Einstein’s equation and then Friedman’s equations for $f(T)$ gravity, we wrote down the continuity equations separately for the above three components. Also, we considered an interaction between the dark components of the universe, such that energy flows from dark matter to dark energy. In what follows, we considered the BEC to be a bosonic dilute gas material that, if cooled to a very low temperature, reaches the quantum ground state. Therefore, we obtained the EoS of BEC as a dark matter as $p_{dm}=\omega_{dm}\rho_{dm}^{2}$. On the other hand, we corresponded the energy density of $f(T)$ gravity with the energy density of new agegraphic. Next, we took the power-law cosmology for the scale factor, which the Hubble parameter has been obtained in terms of the redshift parameter. Then, we plotted the variation of $\rho_{de}$, $p_{de}$, and $\omega_{de}$ in terms of the redshift parameter which showed that the universe is undergoing accelerated expansion. Also, we explored stability analysis of the current model by the sound speed parameter, which the acquired results represented the universe is in a non-stability phase at the late universe. Finally, we test our model by calculating the present values of the density parameters of $\Omega_{bm_{0}}$, $\Omega_{dm_{0}}$, and $\Omega_{de_{0}}$, which indicate the universe is consistent with the observational data.

### v2

@@ IV NA dark energy and BEC dark matter models
where $\dot{\tau}=\frac{1}{a}$. Therefore, the energy density of the NA dark energy reads
@@
where $n^{2}$ is constant.
On the other hand, we study the nature of dark matter by the regime of BEC. BEC is a state of matter in which the dilute Bose gas is cooled to such a low temperature that a phase transition occurs at this point and the macroscopic quantum phenomenon occurs as a quantum minimum state. Since dark matter has a significant contribution to the evolution of the universe, it leads us to adopt the BEC approach instead of dark matter. This means that dark matter is a type of BEC boson gas that dominates a part of the universe. Therefore, we consider the number density of particles by the Bose-Einstein statistics, where these particles form through the decoupling of the remaining plasma in the early universe. However, the energy density of dark matter is equal to the mass of dark matter multiplied by the number density of particles, and dark matter pressure is defined by Bose-Einstein statistics in a sphere with a radius of particles momentum (refer to [92, 93, 94, 95, 96] for more details). In that case, by the generalized Gross-Pitaevskii equation [97, 98], we write down the dark matter pressure in terms of the energy density of dark matter as
@@
where $\omega_{dm}=2\pi l_{a}/m_{dm}^{3}$, in which $l_{a}$ is the s-wave scattering length which here is introduced as dark matter EoS. By inserting Eq. (19) into Eq. (12b), we have
@@
where by inserting $a_{0}=1$ ($a_{0}$ is the current scale factor) we obtain the current $\rho_{dm_{0}}$ as
@@
V Reconstruction and correspondence of the model
In this section, we study the present model based on reconstructing and correspondence methods. In that case, we reconstruct the Friedmann equation in terms of redshift parameter, and then we correspond the NA dark energy with $f(T)$ gravity. For this purpose, we write down the transformation between the scale factor and redshift parameter $z$ as $a(t)=a_{0}/(1+z)$ in which $a_{0}=1$ is the present scale factor, also differential form is as
@@ V Reconstruction and correspondence of the model
Now, in order to motivate the reconstruction of the function $f(T)$, we provide the following scientific explanation to proceed. We note that the modified Friedmann equations in the context of $f(T)$ gravity contain several degrees of freedom, including the energy density and pressure of the distinct components, the Hubble parameter, and the arbitrary function $f(T)$. Without additional constraints, this system of differential equations admits an infinite number of solutions. Therefore, in this work, we use the scale factor $a(t)$, the BEC model for dark matter, and the NA dark energy to reconstruct the function $f(T)$ along with physical motivation and consistency with observational data. These components act as guiding principles to reduce the degeneracy of the solutions and allow us to express the function $f(T)$ explicitly in terms of the torsion scalar $T$. This reconstruction not only increases the predictive power of the model, but also provides deeper insight into the interplay between geometry and quantum matter in the evolution of the universe.
@@
where $t_{0}$ is the present age of the universe, and $m$ is dimensionless positive coefficient. The choice of a power-law scale factor is motivated by both theoretical and observational considerations. Power-law cosmology has been widely used in inflationary models, scalar field dynamics, and modified gravity scenarios due to their analytical tractability and ability to capture key features of cosmic evolution. Observational studies have shown that power-law behavior can approximate the expansion history in certain redshift ranges. By adopting this form, we reduce the arbitrariness in selecting a functional form for $f(T)$ and instead reconstruct it based on physically motivated components and observational constraints. It is important to note that the best-fit procedure primarily constrains the expansion history via the scale factor. This constraint indirectly informs the reconstruction of $f(T)$, which is derived from the modified Friedmann equations and the continuity relations. Therefore, while the fit does not directly test the gravity model, it provides a consistent framework for extracting $f(T)$ from observationally motivated dynamics. In that case, the Hubble’s parameter and consequently the age of the universe will be:
@@
We now analyze the present model using observational constraints with Hubble parameter data against redshift. Observations of the expansion rate of the universe, especially Hubble parameter data $H(z)$ at different redshifts, play a fundamental role in determining and testing cosmological models. Here, relying on experimental data from galaxy observations, cluster convergence, and other independent sources, we extract statistical constraints on theoretical models using rigorous statistical tools. To examine the agreement of the present model with the experimental data, the technique of minimizing the function $\chi^{2}$ is used, which is defined as follows:
$\chi^{2}=\sum\limits_{i=1}^{N}\frac{[H_{th}(z_{i};\overrightarrow{\theta})-H_{obs}(z_{i})]^{2}}{\sigma_{i}^{2}}.$ (27)
[…]
Table 1: The Hubble parameter data set in terms of redshift parameter and their uncertainty values, units of $H(z)$ and $\sigma_{H}$ are $km~s^{-1}~Mpc^{-1}$.
No. Redshift H(z) $\sigma_{H}$ Ref. No. Redshift H(z) $\sigma_{H}$ Ref.
1. 0.07 69.0 19.6 [100] 28. 0.510 90.4 1.9 [108]
2. 0.09 69.0 12.0 [101] 29. 0.52 94.35 2.64 [106]
3. 0.12 68.6 26.2 [100] 30. 0.56 93.34 2.3 [106]
4. 0.17 83 8 [102] 31. 0.57 92.9 7.855 [111]
5. 0.179 75.0 4.0 [103] 32. 0.59 98.48 3.18 [106]
6. 0.199 75.0 5.0 [103] 33. 0.593 104.0 13.0 [103]
7. 0.200 72.9 29.6 [100] 34. 0.6 87.9 6.1 [112]
8. 0.24 79.69 3.32 [104] 35. 0.610 97.3 2.1 [108]
9. 0.27 77 14 [102] 36. 0.64 98.02 2.98 [106]
10. 0.280 88.8 36.6 [100] 37. 0.680 92.0 8.0 [103]
11. 0.30 81.7 5.0 [105] 38. 0.73 97.3 7 [112]
12. 0.31 78.18 4.74 [106] 39. 0.781 105.0 12 [103]
13. 0.34 83.8 2.96 [104] 40. 0.875 125 17 [103]
14. 0.35 82.7 9.1 [107] 41. 0.880 90.0 40.0 [113]
15. 0.352 83 14 [103] 42. 0.900 117 23.0 [102]
16. 0.36 79.94 3.38 [106] 43. 1.037 154 20 [103]
17. 0.38 81.5 1.9 [108] 44. 1.300 168 17 [102]
18. 0.3802 83.0 13.5 [109] 45. 1.363 160 33.6 [114]
19. 0.40 82.04 2.03 [106] 46. 1.430 177 18 [102]
20. 0.4004 77 10.2 [109] 47. 1.530 140 14 [102]
21. 0.4247 87.1 11.2 [109] 48. 1.750 202 40 [102]
22. 0.43 86.45 3.27 [104] 49. 1.965 186.5 50.4 [114]
23. 0.44 84.81 1.83 [106] 50. 2.30 224 8.6 [115]
24. 0.4497 92.8 12.9 [109] 51. 2.33 224 8 [116]
25. 0.470 89 34 [110] 52. 2.340 222 7 [117]
26. 0.4783 80.9 9.0 [109] 53. 2.360 226 8 [118]
27. 0.48 87.79 2.03 [106]
Figure 1: the variations of $H(z)$ in terms of the redshift parameter.
In what follows, we consider that the source of dark energy arises from the $f(T)$ gravity and the NA. Therefore, we correspond Eq. (18) with energy density of dark energy $\rho_{de}$ and obtains as
@@
Now, by entering equations (13), (22), (26), and (28)in Eq. (14a), we can obtain the form of the function $f(T(z))$ as follows:
$f(z)=\frac{2\kappa^{2}\rho_{bm_{0}}}{1-3m(1+\omega_{bm})}(1+z)^{3(1+\omega_{bm})}+\frac{6n^{2}H_{0}^{2}(1-m)^{2}}{m^{2}(2m-1)}(1+z)^{\frac{2(1-m)}{m}}$
$+\frac{2\kappa^{2}b^{2}\rho_{dm_{0}}}{\omega_{dm}}\,_{2}F_{1}\left(1,\frac{1}{3m(1-b^{2})};1+\frac{1}{3m(1-b^{2})};\frac{(1-b^{2})}{\omega_{dm}}(1+z)^{-3(1-b^{2})}\right).$ (30)
The relationship obtained for $f(z)$ is obtained from the modified Friedmann equation in the gravitational framework $f(T)$. This function is explicitly dependent on the redshift $z$ and consists of three independent parts, each of which is related to one of the components of the universe:
The first term: This part is related to the observable baryonic matter, which is characterized by the EoS $\omega_{bm}$ and the initial density $\rho_{bm0}$. Its behavior corresponds to the power $(1+z)^{3}$ in the uncompressed state (dust).
The second term: This part comes from the NA dark energy model, which depends on the parameters $n$, $m$, and $H_{0}$. Its nonlinear power represents the dynamical behavior of dark energy in the cosmological NA framework.
The third term: This part is related to the dark matter of the BEC type coupled with the interaction term $Q$. The presence of the hypergeometric function $_{2}F_{1}$ indicates an analytical solution of the continuity equation with interaction and has a complex dependence on the parameters $b$, $\omega_{dm}$, and $\rho_{dm_{0}}$.
The importance of this explicit relation for $f(z)$ allows us to extract the function $f(T)$ in a reconstructed form from observational data and physical models, and to analyze the dynamical behavior of the universe without assuming an arbitrary form for $f(T)$, but rather using the physical components of the model, and finally, to compare the model with Hubble data and other observations. In that case, using Eqs. (6) and (26), which leads to $1+z=\left(\frac{-T}{6H_{0}^{2}}\right)^{\frac{m}{2}}$, we convert $f(z)$ to the function $f(T)$ as follows:
$f(T)=\frac{2\kappa^{2}\rho_{bm_{0}}}{1-3m(1+\omega_{bm})}\left(\frac{-T}{6H_{0}^{2}}\right)^{\frac{3m(1+\omega_{bm})}{2}}+\frac{6n^{2}H_{0}^{2}(1-m)^{2}}{m^{2}(2m-1)}\left(\frac{-T}{6H_{0}^{2}}\right)^{1-m}$
$+\frac{2\kappa^{2}b^{2}\rho_{dm_{0}}}{\omega_{dm}}\,_{2}F_{1}\left(1,\frac{1}{3m(1-b^{2})};1+\frac{1}{3m(1-b^{2})};\frac{(1-b^{2})}{\omega_{dm}}\left(\frac{-T}{6H_{0}^{2}}\right)^{-\frac{3m(1-b^{2})}{2}}\right).$ (31)
In what follows, we can obtain the dark energy pressure $p_{de}$ by inserting Eq. (28) into the continuity equation (12c), and then we obtain the varition of $\rho_{de}$, $p_{de}$, and $\omega_{de}$ for dark energy component in terms of redshift parameter as shown in Figs. 2 and 3. The choice of values of free parameters in cosmological models plays a crucial role in their physical validity and agreement with observational data. The combination of these parameters is chosen in such a way that the reconstructed function $f(T)$ is not only consistent with experimental data, but also has theoretically acceptable behavior at different epochs of the universe, including the transition from the matter era to the accelerated era. Therefore, the free parameters are selected as $n=18$, $b=0.807$, $\omega_{dm}=2.5$, $\rho_{dm_{0}}=4873$, $\omega_{bm}=1$, and $\rho_{bm_{0}}=545$ which $n$, $b$, $\omega_{dm}$, and $\omega_{bm}$ are dimensionless quantities, and $\rho_{dm_{0}}$ and $\rho_{bm_{0}}$ are dimension $M^{4}$ in Planck units. We can see in Figs. 2 that the value of energy density decrease from a high positive value in the early universe to a lower positive value in the late universe ($z=0$), also the value of pressure decreases from a high positive value in the early universe to a lower negative value in the late universe ($z=0$).
@@
We note that the EoS parameter is a very important parameter so that can understand at a glance how the dynamics of the universe are evolving from the early time to the late time. Therefore, we notice that the matter period and the accelerated pase period for the universe can be expressed as $0<\omega<\frac{1}{3}$ and $1<\omega<-\frac{1}{3}$, respectively. Hence, Fig. 3 shows us that the universe evolves from the matter-dominated era to the accelerated expansion era in which the value of EoS is about 0.71-0.71 in the present time ($z=0$), indicating that the universe enters a thermodynamically unstable phase in the late-time evolution.
@@
In what follows, we explore the stability of the model by using the sound speed parameter which is derived from thermodynamic considerations. For this reason, we consider the universe to be an adiabatic thermodynamic system where no heat or mass is transferred out, meaning that the perturbation has zero entropy. Therefore, the sound speed parameter $c_{s}^{2}$ is defined in the following form
@@
where the index $z$ indicates derivative with respect to redshift parameter. We plotted the sound speed parameter as a function of redshift parameter as shown in Fig. 4. We note that the conditions of larger and smaller sound speed parameters, i.e., $c_{s}^{2}>0$ and $c_{s}^{2}<0$, indicate stability and instability, respectively. In that case, Fig. 4 displays that the universe begins from a stability phase and then at the threshold of the accelerated phase enters a non-stability phase in late time. Therefore, this variation represents that the energy density of dark energy in the late universe is not in a controlled growth.
Figure 4: The graph of the sound speed (dimensional quantity) in terms of redshift parameter with the same values of $m=0.956$, $n=18$, $b=0.807$, $\omega_{dm}=2.5$, $\rho_{dm_{0}}=4873$, $\omega_{bm}=1$, and $\rho_{bm_{0}}=545$.
@@
VI Conclusion
In this paper, we studied the cosmic evolution of the universe in the $f(T)$ modified gravity framework with a smooth Friedmann-Robertson-Walker (FRW) background. We considered the universe to consist of three components: baryonic matter, dark matter, and dark energy, where dark matter was replaced by the BEC model and dark energy by the NA dark energy model. After deriving the Einstein field equations and Friedmann equations in the $f(T)$ background, we wrote the continuity equations for each component of the universe separately and considered an interaction between dark matter and dark energy in which energy is transferred from dark matter to dark energy.
For the dynamical analysis, we chose power-law cosmology for the scale factor and reconstructed the Hubble parameter in terms of the redshift parameter $z$. Using the 53-point Hubble parameter data, the optimal value of the power parameter $m=0.956$ was obtained, which led to the calculation of the current age of the universe as $t_{0}=13.87~Gyr$. This value is fully consistent with recent observational data.
Next, using the BEC model for dark matter and the NA dark energy model for dark energy, we obtained the function $f(T)$ explicitly and reconstructed in terms of the torsion scalar $T$. This function consists of three independent parts:
- The first part: related to baryonic matter with a strong dependence on $(1+z)^{3}$,
- The second part: related to the dark energy of the new sensor with a nonlinear dependence on the parameters $n$, $m$ and $H_{0}$,
- The third part: related to BEC dark matter type with the presence of the interaction term and the hypergeometric function $_{2}F_{1}$ which represents the analytical solution of the continuity equation with interaction.
Plots of energy density $\rho_{\text{de}}$, pressure $p_{\text{de}}$, and EoS parameter $w_{\text{de}}$ versus redshift showed that the universe moves from a matter-centered phase to an accelerated expansion, and the value $w_{\text{de}}\approx-0.71$ was obtained at the present time ($z=0$). This value indicates that the universe enters an accelerated and unstable phase at later times.
To investigate thermodynamic stability, we calculated the speed of sound parameter $c_{s}^{2}$. The corresponding plot showed that the universe transitions from a stable phase at early times to an unstable phase at later times, which is consistent with the behavior of dark energy in dynamical models.
Finally, by calculating the current density parameters for baryonic matter, dark matter, and dark energy as $\Omega_{bm_{0}}=0.04$, $\Omega_{dm_{0}}=0.27$, and $\Omega_{de_{0}}=0.69$, we show that the proposed model is in good agreement with recent observational data, including CMB results and large-scale structure.
It is worth noting that in the present model, the roles of these three components are physically distinct, although they are connected through the Friedmann background. The function $f(T)$ determines the geometric evolution of the universe through torsion and provides the gravitational framework. The BEC dark matter, through its nonlinear equation of state, imposes quantum corrections that mainly affect the matter-dominated period and structure formation. The NA dark energy component, in turn, drives the finite-time acceleration through the isotropic time dependence of its energy density. Hence, while these three components interact dynamically, their individual actions remain separate and contribute differently to cosmic evolution.
In addition to the numerical and analytical results obtained, it can be emphasized that the main objectives of the research have been fully realized. The proposed model was able to provide a coherent picture of the evolution of the universe by combining the three theoretical frameworks of $f(T)$ gravity, BEC dark matter, and NA dark energy, which is consistent with observational data. The reconstruction of the $f(T)$ function based on the physical components of the model and the $H(z)$ data showed that the proposed model is able to explain the cosmic acceleration without the need for a cosmological constant. Also, the analysis of thermodynamic stability and the behavior of the equation of state parameter confirmed that the transition from the matter-dominated phase to the acceleration-dominated phase is consistent with theoretical predictions. Therefore, the results of this research indicate that the assumptions and objectives set at the beginning of the work, including the geometric-quantum description of the universe and compatibility with observations, have been achieved quantitatively and qualitatively. In the future, this model can be extended by analyzing perturbations, structure formation, and CMB and BAO data to further examine its validity.